OpenStudy (anonymous):

In a classroom, 5 out of 13 students surveyed said they had two or more siblings. If the entire school has 520 students, predict the number of students that have two or more siblings. Write a proportion to model the problem.

OpenStudy (johnweldon1993):

Well we can write the proportion as \[\large \frac{5}{13} = \frac{x}{520}\]

OpenStudy (johnweldon1993):

Then to solve it...you would cross multiply \[\large 2600 = 13x\] And then you would divide both sides by 13 to solve for 'x' \[\large x = \frac{2600}{13} = \space ? \]

OpenStudy (anonymous):

thanks i have one more.... In a classroom, 11 out of 17 students surveyed say they had two or more siblings. If the entire school has 850 students, predict the number of students that have two or more siblings.

OpenStudy (johnweldon1993):

So same thing there... \[\large \frac{11}{17} = \frac{x}{850}\] \[\large 9350 = 17x\] \[\large x = \frac{9350}{17} = \space ? \]

OpenStudy (anonymous):

oh.. ok thanks.. i thought so!!! i am just checking my answers

OpenStudy (johnweldon1993):

No problem :)